Please help. I am lost.
Manufacturing shirts. If a manufacturer charges p dollars each for rugby shirts, then he expects to sell 2000 – 100p shirts per week. What polynomial represents the total revenue expected for a week? How many shirts will be sold if the manufacturer charges $20 each for the shirts? Find the total revenue when the shirts are sold for $20 each. Use the bar graph to determine the price that will give the maximum total revenue.

Answer by stanbon(75887)

If a manufacturer charges p dollars each for rugby shirts, then he expects to sell 2000 – 100p shirts per week.
-------------
What polynomial represents the total revenue expected for a week?
Revenue = (price)*(quantity) = p(2000-100p) = 2000p-100p^2
--------------
How many shirts will be sold if the manufacturer charges $20 each for the shirts?
# = 1000-100(20) = 0
No shirts will be sold
-----------------------
Find the total revenue when the shirts are sold for $20 each. Use the bar graph to determine the price that will give the maximum total revenue.
-------
I don't have the bar graph but maximum revenue occurs when
p = -b/2a = -2000/(2*-100) = $10
-----------------
Cheers,
Stan H.